(2cos40 o - cos20 0 ) / sin20 0 Can you help me ,what is the result

(2cos40o - cos200) / sin200 
Can you help me ,what is the result 

Grade:12th pass

1 Answers

Deepak Kumar Shringi
5 years ago
2*cos40 - cos20 = cos40 + (cos40 - cos20) = cos40 - 2*sin((40+20)/2)*sin((40-20)/2) = cos40 - 2*sin30*sin10 = sin(90-40) - sin10 = sin50 - sin10 = 2*sin((50-10)/2)*cos((50+10)/2)=2*sin20*√3/2=√3*sin20
√3*sin20 / sin20 = √3